If we multiply both sides of equation 21.10 by Bav, we get 8pv (21.33) Notice that since the metric with upstairs indices is the matrix inverse of the ver- sion with downstairs indices, gavg㎛-δμμ 4. Exercise 21.3.1. Why is this equal to 4? You can use this and the definitions of R and T to show that-R + 4A = KT. Exercise 21.3.2. Verify this.

Definitions:

media%2F086%2F086c6afe-e912-4987-a2a3-fb


media%2Fbe8%2Fbe86d63c-2adf-4bff-bc89-dc



media%2F0c8%2F0c8f947a-362a-409e-949a-45

media%2Fc54%2Fc545fb6f-a450-437d-8d16-65
Answer #1

1) In the equation

\small g_{\mu\nu}g^{\mu\nu}=\delta_\mu^{\mu}=4

4 is the dimension of the space-time. This means the metric tensor \small g_{\mu\nu} is a 4-dimensional square matrix.

2) In general relativity, a scalar can be defined using metric tensor and a rank 2 tensor as follows

\small g_{\mu\nu} T^{\mu \nu}=T

T is called trace of the energy-momentum tensor \small T^{\mu \nu}

similarly

\small g_{\mu\nu} R^{\mu \nu}=R

R is the Ricci scalar which is obtained taking the trace of the Ricci tensor

Hence with these definitions, the equation 21.33 becomes

\small R-\frac{1}{2}\times 4 \times R+\Lambda \times 4=\kappa T

This can be written as

\small R(1-2)+4\Lambda =\kappa T

which gives

\small -R+4\Lambda =\kappa T